II. Fill out the loan amortization schedule provided in the solution template for the first 5 loan payments. What do you notice about the composition of the payment amount?
1
Expert's answer
2020-06-18T18:21:20-0400

Beginning balance computation

Amount = $326,000

Deposit = $75,000

Balance = Amount – Deposit

Balance = $326,000 - $75,000 = $ 251,000

Beginning balance 1 = $251,000.00

Beginning balance 2 = $250,862.90

Beginning Balance 3 = $ 250,724.77

Beginning Balance 4 = $250,445.39



Interest Payment computation

Interest rate = 0.75% per month.

Interest = Beginning balance * Interest rate per month

Interest 1= 251,000 (0.75%) = $1882.50

Interest 2 = $250,862.90 * 0.75% = $ 1,881.47

Interest 3 = $ 250,724.77 * 0.75% = $ 1,880.44   

Interest 4 = $250,585.60 * 0.75% =$ 1,879.39

Interest 5 = $ 250,445.39 * 0.75% =$ 1,878.34

 


Principal Payment computation

Principal = PMT / (1.0075)³⁶⁰

Principal = 2019.6 / 14.730576 = $137.10

Principal = 2019.6 / 14.730576 = $ 138.13

Principal = 2019.6 / 14.730576 = $ 139.17

Principal = 2019.6 / 14.730576 = $ 140.21

Principal = 2019.6 / 14.730576 = $ 141.26

  

Ending Balance

Ending Balance = Beginning balance – Principal

Ending Balance 1= $251,000 – $137.10 = $ 250,862.90

Ending Balance 2= $250,862.90 – 138.13 = $ 250,724.77   

Ending Balance 3= $250,724.77 – $ 139.17 = $250,585.60

Ending Balance 4= $250,724.77 – $ 140.21 = $ 250,445.39

Ending Balance 5= $250,445.39 – $ 141.26 = $ 250,304.13

 

From the loan amortization schedule, it can be noted that the principal increases at a uniform rate/amount of $1.03 and $1.05 over the 5 year period. The principal incremental rate is spread out across the 5 year period.